OpenStudy (anonymous):

5.2 over 3.5 = y over 7.5 I can't figure out y

OpenStudy (anonymous):

give me a min

OpenStudy (campbell_st):

\[\frac{5.2}{3.5} = \frac{y}{7.5}\] multiply both sides of the equation by 7.5 \[\frac{5.2 \times 7.5}{3.5} = y\] just calculate the answer

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no use butterfly method it is easyer

OpenStudy (anonymous):

5.2 times 7.5 divided by 3.5 and there is y

OpenStudy (anonymous):

Thank you too

OpenStudy (anonymous):

np

OpenStudy (anonymous):

y = 11.3 or 11.14